OpenStudy (meilendurcer):

The functions f(x) and g(x) are described using the following equation and table. f(x) = 4(1.02)x x g(x) -1 -4 0 6 1 8 2 10 Which statement best compares the y-intercepts of f(x) and g(x)? A. The y-intercept of f(x) is equal to the y-intercept of g(x). B. The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). C. The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). D. The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (anonymous):

Is the first row in the table typed incorrectly? I think it should be g(x) = 4 (not -4) for x = -1, because then g(x) will be a linear function

OpenStudy (meilendurcer):

@navk yes its typed correctly.

OpenStudy (anonymous):

if you plot the points for g(x), you can see that its y-coordinate is 6

OpenStudy (anonymous):

f(x) has a y-intercept at x = 0 \[f(0) = 4(1.02))^0 = 4(1) = 4\]

OpenStudy (meilendurcer):

would it be A or D? i'm kind of stuck.

OpenStudy (anonymous):

y-intercept of g(x) = 6 y-intercept of f(x) = 4 so D

OpenStudy (meilendurcer):

ok i see thank you. Can you help with a few more if possible??

OpenStudy (anonymous):

np

OpenStudy (meilendurcer):

A. 7 1/2 B. 7 8/15 C. 7 5/3 D. 7 2/15

OpenStudy (meilendurcer):

i have no idea what the answer would be.

OpenStudy (anonymous):

it is cube root of 7 over fifth root of 7, which can be written as \[\frac{7^{1/3}}{7^{1/5}}\]then subtrat the exponents 1/3-1/5=2/15

OpenStudy (meilendurcer):

thanks! i understand.

OpenStudy (anonymous):

great!

OpenStudy (meilendurcer):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

To get the value of x, you need to get rid of y. Any ideas how to do that ?

OpenStudy (meilendurcer):

not exactly

OpenStudy (meilendurcer):

subtract 5 from 3??

OpenStudy (anonymous):

no first you have to multiply the first equation by 3 so that in both equations the coefficient of y becomes the same, that is, 3

OpenStudy (anonymous):

So you get 3*{x − y = −3} = 3x - 3y = -9 Now add this and the second equation to get {3x - 3y} + {x + 3y} = -9 + 5 4x = -4 x = -1

OpenStudy (meilendurcer):

ahh.... ok